OpenStudy (anonymous):

\[\frac{1}{a^3}+\frac{1}{b^3}=?\] if a+b= x and ab=y

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

take LCM, its done

OpenStudy (anonymous):

\[\frac{a^3+b^3}{a^3b^3}\]

OpenStudy (anonymous):

af

OpenStudy (anonymous):

after that ?

OpenStudy (anonymous):

wait one mint

OpenStudy (anonymous):

x(b^2 -y +a^2)/a^3b^3

OpenStudy (anonymous):

a^2+b^2=a^2+b^2+2ab-2ab=(a+b)^2-2ab

OpenStudy (anonymous):

a^2b^2=(ab)^2

OpenStudy (anonymous):

and put the given values

OpenStudy (anonymous):

am not able to solve further please show the full process

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

[a^2+b^2]/[a^2b^2] =[a^2+b^2+2ab-2ab]/[(ab)^2] =[(a+b)^2-2ab]/[(ab)^2] =[x^2-2y]/[y]^2

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

that is okay but it should be (ab)^3 or y^3 in the denominator but i understood the process thanks

OpenStudy (anonymous):

u simplify, whatever is it

OpenStudy (anonymous):

no no sorry i think the whole process went wrong zakaullah ,,,, i would request u to correct it and post it here please so that i can understand it !!!!!!!!!!!!!!!!

OpenStudy (anonymous):

its correct, just think about it a little bit

OpenStudy (anonymous):

\[\frac{a^3+b^3}{a^3b^3}\] \[\frac{(a+b)(a^2+b^2+ab)}{a^3b^3}\] now ?

OpenStudy (anonymous):

here it looks like a^2, sorry for that

OpenStudy (anonymous):

it's \(\frac{1}{a^3} + \frac{1}{b^3}\) right?

OpenStudy (anonymous):

k no problem is it right now : \[\frac{a^3+b^3}{a^3b^3}\] \[\frac{(a+b)(a^2+b^2+ab)}{a^3b^3}\] \[\frac{(a+b){(a+b)^2-ab)}}{y^3}\] \[\frac{x(x^2-ab)}{y^3}\] \[\frac{x^3-xy}{y^3}\]

OpenStudy (anonymous):

in the above [a^2+b^2+ab] =[a^2+b^2+2ab-ab] =(a+b)^2-ab now try

OpenStudy (anonymous):

m i right ?

OpenStudy (anonymous):

yes 100%

OpenStudy (anonymous):

it is showing the answer as \[\frac{x^3-3xy}{y^3}\]

OpenStudy (zarkon):

you didn't factor the sum of cubes correctly

OpenStudy (anonymous):

why zarkon sir ?

OpenStudy (zarkon):

\[a^3+b^3=(a+b)(a^2+b^2-ab)\]

OpenStudy (anonymous):

k no problem is it right now : \[\frac{a^3+b^3}{a^3b^3}\] \[\frac{(a+b)(a^2+b^2-ab)}{a^3b^3}\] \[\frac{(a+b){(a+b)^2+3ab)}}{y^3}\] \[\frac{x(x^2+3ab)}{y^3}\] \[\frac{x^3-3xy}{y^3}\]

OpenStudy (anonymous):

now it is right ?

OpenStudy (zarkon):

the work is not totally correct

OpenStudy (anonymous):

yes it may be some where wrong since i didn't correct it perfectly ok zarkon sir thanks very much for your help ... greatful to you sir !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

hav to go sir bye

OpenStudy (zarkon):

\[\frac{(a+b)(a^2+b^2-ab)}{a^3b^3}=\frac{(a+b)((a+b)^2-3ab)}{a^3b^3}=\cdots\]

OpenStudy (anonymous):

yes sir i got it i put + and in answer i put - . thanks again

OpenStudy (zarkon):

good